Rob spent $64 to make an advanced purchase of 10 movie tickets for himself and his friends. If adult tickets cost $8.50 and senior discount tickets cost $5, determine the number and type of tickets purchased.
Answer by ewatrrr(24785)


Hi
Let x represent number of Adult tickets (most expensive)
$8.50x + $5(10-x) = $64
x = 14/3.5 = 4, the number of Adult tickets. 6 Student tickets 10-4
And
4($8.50) + 6($5) = $64 CHECKS!